OpenStudy (anonymous):

I really need help understanding this question please. I believe its dealing with inequalities. You are going to invest 25,000, part at 14% and part at 16%. What is the most that can be invested at 14% to make at least 3600 interest per year?

OpenStudy (anonymous):

Let amount invested at 14% be "x". Interest earned in one year on this investment is 0.14x dollars. Amount invested at 16% is "25000-x". Interest earned in one year on this investment is 0.16(25000-x) = 4000-0.16x dollars. INEQUALITY: interest + interest >=3600 dollars 0.14x+4000-0.16x >=3600 -0.02x>=-400 x<=$20,000 (amount to be invested at 14%)

OpenStudy (anonymous):

where did the 4000 come from?

OpenStudy (anonymous):

@ILoveMath3006

OpenStudy (anonymous):

.14x +.16(25000-x) >= 3600 .14x +4000 -.16x >=3600 -.02x>=-400 x<=20000 $20 000 Well let me put it like this

OpenStudy (anonymous):

As you see, the 4000 come from .16(25000-x)

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Np